OpenStudy (anonymous):

find \[f ^{-1}\] (x) f(x) = sqrt[3]{x-7}

jimthompson5910 (jim_thompson5910):

First you replace f(x) with y. Then you swap x and y. Afterwards you solve for y to get the inverse \[\Large f(x) = \sqrt[3]{x-7}\] \[\Large y = \sqrt[3]{x-7}\] \[\Large x = \sqrt[3]{y-7}\] \[\Large x^3 = y-7\] I'll let you finish

OpenStudy (anonymous):

Im a little lost still would it be \[y=x ^{3}+7\]

jimthompson5910 (jim_thompson5910):

you are correct, the last step would be to replace y with the inverse function notation symbol to get \[\Large f^{-1}(x) = x^3 + 7\] but you pretty much have it

OpenStudy (anonymous):

oh awesome it just didn't seem right this math is killing me

jimthompson5910 (jim_thompson5910):

just gotta trust yourself a bit more and keep practicing it'll get better

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw